# Deep Reinforcement Learning: A Convex Optimization Approach

### 07 Mar 2024 v2

@@ 2 Episodic Deep Reinforcement Learning with Convex Optimization
Further, let $\alpha_{k}$ be either a positive real constant, $\alpha_{k}=\alpha<1$ for all $k\in\mathbb{N}$, or such that $\alpha_{k}=1/k^{q}$ with $q>1$. If
@@
$\|w_{k}-w^{\star}\|\leq C\cdot\frac{\rho}{T}$
@@ 4 Proof of Theorem 1
$11-\alpha_{k}+\alpha_{k}\frac{3\gamma^{2}\beta^{2}}{2\lambda^{2}T}<$ 1 .
For the case where $\alpha_{k}=\alpha$, for all $k\in\mathbb{N}$, we see that the inequality in (10) implies that $w_{k}-w^{\star}$ converges and that
$\|w-w^{\star}\|\leq\frac{1}{1-\sqrt{\frac{3\gamma^{2}\beta^{2}}{2\lambda^{2}T}}}\frac{\rho f_{\textup{max}}}{\lambda T}\|w^{\star}\|$
Similarly, for $\alpha_{k}=\frac{1}{k^{q}}$, $\|w_{k}-w^{\star}\|$ is the state $\Delta_{k}$ of a stable dynamical system given by
$\Delta_{k+1}=a_{k}\cdot\Delta_{k}+\alpha_{k}\frac{\rho f_{\textup{max}}}{\lambda T}\|w^{\star}\|$
where
$a_{k}=1-\alpha_{k}+\alpha_{k}\frac{3\gamma^{2}\beta^{2}}{2\lambda^{2}T}=1-\frac{1}{k^{q}}\left(1-\frac{3\gamma^{2}\beta^{2}}{2\lambda^{2}T}\right)<1.$
Introduce
$b=1-\frac{3\gamma^{2}\beta^{2}}{2\lambda^{2}T}$
Now for any $K$, we have that
$\Delta_{K}\leq\sum_{k=1}^{K}\left(1-\frac{b}{k^{q}}\right)^{k}\frac{1}{k^{q}}\frac{\rho f_{\textup{max}}}{\lambda T}\|w^{\star}\|$ (11)
@@
$\sum_{k=1}^{\infty}\left(1-\frac{b}{k^{q}}\right)^{k}\frac{1}{k^{q}}<\infty$
for $q>1$, we may set
$C=\sum_{k=1}^{\infty}\left(1-\frac{b}{k^{q}}\right)^{k}\frac{1}{k^{q}}\frac{f_{\textup{max}}}{\lambda}\|w^{\star}\|$
and thus, obtain the inequality
$\|w_{k}-w^{\star}\|=\Delta_{k}\leq C\cdot\frac{\rho}{T}$
for some positive real constant $C$. This completes the proof.

### 20 Mar 2024 v3

@@ 2 Episodic Deep Reinforcement Learning with Convex Optimization
Further, let $\alpha_{k}$ be either a positive real constant, $\alpha_{k}=\alpha<1$ for all $k\in\mathbb{N}$, or such that $\alpha_{k}=1/k$. If
@@
$\|w-w^{\star}\|\leq C\cdot\frac{\rho}{T}$
@@ 4 Proof of Theorem 1
Introduce
$\mu=1-\sqrt{\frac{3\gamma^{2}\beta^{2}}{2\lambda^{2}T}}<1$
@@
$11-\mu\alpha_{k}<$ 1 .
For the case where $\alpha_{k}=\alpha$, for all $k\in\mathbb{N}$, we see that the inequality in (10) implies that $w_{k}-w^{\star}$ converges to some $w-w\star$ where
$\|w-w^{\star}\|\leq\frac{\mu\alpha}{1-\mu\alpha}\frac{\rho f_{\textup{max}}}{\lambda T}\|w^{\star}\|$
Similarly, for $\alpha_{k}=\frac{1}{k}$, we have that $\|w_{k}-w^{\star}\|$ is the state $\Delta_{k}$ of a stable dynamical system given by
$\Delta_{k+1}\leq(1-\mu\alpha_{k})\cdot\Delta_{k}+\alpha_{k}\frac{\rho f_{\textup{max}}}{\lambda T}\|w^{\star}\|$
More explicitely, we have that
$\Delta_{K}\leq\prod_{k=1}^{K}(1-\mu\alpha_{k})\Delta_{0}+\sum_{k=1}^{K}\prod_{i=k+1}^{K}(1-\mu\alpha_{i})\alpha_{k}\frac{\rho f_{\textup{max}}}{\lambda T}\|w^{\star}\|$ (11)
Note that
$\prod_{k=1}^{K}(1-\mu\alpha_{k})$ $\leq\prod_{k=1}^{K}\exp(-\mu\alpha_{k})$ (12)
$=\exp(-\mu\sum_{k=1}^{n}K\alpha_{k})$
$\leq\exp(-\mu\ln{K})$
$=\frac{1}{K^{\mu}}$
which goes to zero as $K\rightarrow\infty$. Similarly, for $k>0$, we have
$\prod_{i=k+1}^{K}(1-\mu\alpha_{i})$ $\leq\prod_{i=k+1}^{K}\exp(-\mu\alpha_{i})$ (13)
$=\exp\left(-\mu\sum_{i=k+1}^{K}\alpha_{i}\right)$
$\leq\exp\left(-\mu(\ln{n}-\ln k-1)\right)$
$=\left(\frac{ek}{K}\right)^{\mu}$
Thus,
$\sum_{k=1}^{K}\prod_{i=k+1}^{K}(1-\mu\alpha_{i})\alpha_{k}$ $\leq\sum_{k=1}^{K}\left(\frac{ek}{K}\right)^{\mu}\frac{1}{k}$ (14)
$=\left(\frac{e}{K}\right)^{\mu}\sum_{k=1}^{K}\left(\frac{1}{k}\right)^{1-\mu}$
[…]
$=\left(\frac{e}{K}\right)^{\mu}\frac{1}{\mu}\left((K+1)^{\mu}-1\right)$
@@
$\lim_{K\rightarrow\infty}\frac{e^{\mu}}{\mu}\frac{1}{K^{\mu}}\left((K+1)^{\mu}-1\right)=\frac{e^{\mu}}{\mu}$
we get the inequality
$\sum_{k=1}^{\infty}\prod_{i=k+1}^{\infty}(1-\mu\alpha_{i})\alpha_{k}$ $\leq\frac{e^{\mu}}{\mu}$ (15)
We conclude that
$\|w-w^{\star}\|=\Delta_{\infty}\leq C\cdot\frac{\rho}{T}$
with
$C=\frac{e^{\mu}}{\mu}\frac{f_{\textup{max}}}{\lambda}\|w^{\star}\|.$
This completes the proof.